A)Determine the following integrals:

(i) "\int" (x-2/x^2)(X+2/x^2)dx

(ii) "\int" e^5x(e^2x/7+3/e^3x)dx

(iii) "\int" 1/(4-√3x)^3 dx

(iv) "\int" π/40(tan x)^3(sec x)^3 dx


B)let f(x)=x^2-2 and g(x)=-|x|,then

(i) sketch the graphs of f and g on the same axes.

(ii) find the area enclosed by f(x)=x^2-2 and g(x)=-|x|.




1
Expert's answer
2021-09-05T17:59:31-0400

"A\\
i)\\
\int [x-\frac{2}{x^2}][x+\frac{2}{x^2}]dx\\
=\int [x^2-\frac{4}{x^4}]dx\\
= \frac{x^3}{3}+\frac{4}{3x^3}+c\\


ii)\\
\int e^{5x}[\frac{e^{2x}}{7}+\frac{3}{e^{3x}}]dx\\
=\int [\frac{e^{7x}}{7}+3{e^{2x}}]dx\\
=\frac{e^{7x}}{49}+\frac{3e^{2x}}{2}+c\\



iii)\\

\int \frac{1}{(4-\sqrt{3x})^3}dx\\

\text{put}\\

4-\sqrt{3x}=t\\

=>\frac{-3}{2\sqrt{3x}}dx=dt\\

\implies dx=\frac{-2}{3}(4-t)dt\\

\int \frac{1}{(4-\sqrt{3x})^3}dx\\

=\frac{-2}{3} \int \frac{4-t}{t^3}dt\\

=\frac{-2}{3} \int (4t^{-3}-t^{-2})dt\\

=\frac{-2}{3}(\frac{-2}{t^{2}}+\frac{1}{3t^3})+c\\

=\frac{-2}{3}(\frac{-2}{(4-\sqrt{3x})

^{2}}+\frac{1}{3(4-\sqrt{3x})^3})+c\\


iv)\\
\int^{π/4}_{0 } (tan x)^3 (sec x )^3dx\\ 
\text{put},\\
secx=t\\
=>tanx \space secx \space
 dx=dt\\
\text{and using }tan^2x=1-sec^2x,
\text{we get}\\
\int^{π/4}_{0 } (tan x)^3 (sec x )^3dx\\
=\int^{π/4}_{0 } (secx)^2(tan x )^2tanx \space secx\space dx\\
=\int^{π/4}_{0} (sec x )^2(1-sec^2x)tanx \space secx\space dx\\
 =\int^{\sqrt2}_1 t^2(1-t^2)dt\\
=\int^{\sqrt2}_1 (t^2-t^4)dt\\
=[\frac{t^3}{3}-\frac{t^5}{5}]^{\sqrt2}_1\\
=-0.32\\




B\\
i)"



blue curve denotes "f=x^2-2"

red curve denotes g=-|x|

(ii)

area of region I i.e., ABC,

"2×\frac{1}{2}×1×1\\
=1"

area of region II i.e., ADC,

"2×\int_{-1}^1 -(x^2-2)dx\\
=2×\int_{-1}^1 (2-x^2)dx\\
=2[2x-\frac{x^3}{3}]_{-1}^1\\
=2[2-\frac{1}{3}+2-\frac{1}{3}]\\
=\frac{20}{3}"

total area of the region bounded i.e., ABCD is 23/3.